### v1

# Data Matters:

Abstract
Accurate predictions of base stations’ traffic load are essential to mobile cellular operators and their users as they support the efficient use of network resources and sustain smart cities and roads. Traditionally, cellular network time-series have been considered for this prediction task. More recently, exogenous factors such as points of presence and other environmental knowledge have been introduced to facilitate cellular traffic forecasting. In this study, we focus on smart roads and explore road traffic measures to model the processes underlying cellular traffic generation with the goal to improve prediction performance. Comprehensive experiments demonstrate that by employing road flow and speed, in addition to cellular network metrics, cellular load prediction errors can be reduced by as much as $56.5\%.$ The code and more detailed results are available on https://github.com/nvassileva/DataMatters.
Index Terms:
data, machine learning, mobile cellular traffic, forecasting.
I Introduction
Accurate mobile cellular traffic predictions are critical to mobile operators as the efficient use of network resources and delivery of services depends largely on them. Smart cities and roads are highly dependent on the availability of mobile cellular infrastructure too. Traditionally, the traffic load at a base station (BS) is predicted based on mobile cellular key performance indicators (KPIs). Recently, with the introduction of deep learning models [1], exogenous factors, in addition to cellular time-series data, are incorporated to further facilitate the forecasting process. Typically considered are points-of-presence such as commercial centers, touristic spots or sport venues. More recent works [2] broaden this contextual modeling to include more environmental factors that characterize the surroundings of each BS. In an effort to capture mobility dynamics the handover rate is harnessed in [3]. Nonetheless, data that captures population dynamics remains mainly unexplored. Our goal is to model better the cellular traffic generation process by employing information about the density of the potential sources of cellular load together with the cellular time-series data. In contrast to the majority of the works, we focus on highways instead of urban scenarios and on short-term forecasting, which is arguably more challenging than mid- and long-term forecasting yet relevant to low-latency and high-reliability services, especially those related to smart traffic and autonomous driving.
Our major contributions can be summarized in:
  • We capture the intricate dynamics of the mobile cellular traffic generation process by combining road traffic metrics that account for the potential sources of cellular communication load with cellular traffic time series data.
  • We develop a methodology for generating cellular traffic volumes based on road traffic.
  • We conduct comprehensive experiments for a real-world highway under a diversity of mobile cellular and road traffic conditions, which complement our first study [4].
The methodology we develop for creating cellular traffic data sets based on road traffic is explained in Section II. A description of the real highway scenario used in the experiments can be found in Section III. The traffic prediction model and the learning model we use for solving it are formulated in Section IV. We discuss the experimental setting and performance results in Section V, and conclude the paper with a summary of the main results and future prospects in Section VI.
II Methodology
Measuring and recording a large variety of different KPIs is intrinsic to any communication network because the measurements are used for monitoring, operating, optimizing and maintaining the network. However, mobile operators rarely share such data because of privacy concerns. One of the very few open access data sets [5] is the Telecom Italia Big Data Challenge data [6]. Therefore, the majority of the studies, which develop machine learning methods for predicting mobile cellular traffic, routinely use it. Commonly, their goal is to predict the number of calls as a measure of the cellular network load. We follow the same approach of using calls.
Similar to the KPIs’ relevance to mobile cellular networks, road traffic measurements are critical to any off-the-shelf traffic management system. In contrast to cellular KPIs, however, traffic measurements are open access. The California Department of Transportation (Caltrans), for instance, is maintaining a Performance Measurement System (PeMS) [7] with real-time traffic data from a large number of individual detectors deployed statewide in the California freeway system. Both real-time and historical measurements are freely available.
We use real-life PeMS data to generate mobile cellular traffic due to the lack of open data sets comprising both network KPIs and road traffic measurements. Before describing our methodology, we summarize the variables we use from PeMS.
II-A Road data
The road metrics employed in this study are vehicular flow—the number of vehicles that pass by a specific location during a time interval—and speed of the flow. In the PeMS data set, the former metric is aggregated over five-minute intervals and the latter is averaged over these intervals.
We make an important assumption that road traffic is measured outside yet in the vicinity of the BS of interest, so that the flow and speed measured in time slot $t$ are observed during time slot $t+1$ in the area covered by this BS.
II-B Mobile cellular data generation
We use the described PeMS variables to emulate the time of arrival of the vehicles per five-minute time slot, the dwell time they spent in the cell and their departure times. These, together with the call duration, are then used to determine if a call is handed over to the next BS on the road or if it is terminated before the vehicle leaves its serving BS.
II-B1 Vehicular flow arrival and departure
The loop detectors of the California transportation system measure all variables in real time yet the measurements are recorded in PeMS as aggregates for the number of vehicles and as averages for the speed per five-minute intervals [^1] The five-minute interval is typical for most transportation systems. as mentioned earlier. Vehicular arrival and departures times at a BS are needed to simulate the call cellular load. We use the conventional Poisson assumption to emulate vehicular arrival times. Some more recent studies [8] validate this hypothesis with real data from highways for sparse—not congested nor experiencing bursts—traffic. Since for the large majority of the data we employ, the speed is around its maximum allowed value (70 mph), suggesting congestion is not habitual, we adopt this assumption too: we emulate the vehicles’ arrival time by exponential inter-arrival times within each five-minute interval. The vehicle’s departure time is determined by its arrival and dwell times.
II-B2 Dwell time
The time spent by a vehicle in a segment of a highway served by a BS, is modeled by the range of the BS and the measured PeMS average speed. We add Gaussian noise with mean zero and standard deviation 0.05 to emulate different speeds.
II-B3 Calls
New call arrivals at a BS follow a Poisson process (see [9] for an overview). Therefore, we model the calls placed by each vehicle by exponential call inter-arrival times. Each vehicle generates calls with a given rate $\lambda$, which is fixed and common for all vehicles. The call arrival process at a BS then is a non-homogeneous Poisson process, with arrival rate given by the sum of the time-varying vehicular arrival rate (different per five-minute interval) and the fixed call arrival rate.
Figure 1: A highway with merging arterial roads and BS coverage.
We do not model the handover arrival process explicitly. The number and rate of handovers are determined by the start and end times of the simulated calls and the vehicle’s departure time from a BS. Hence, the handover process is completely defined by the call duration, cell range, vehicle’s speed and new call generation process. If the call is not terminated before the vehicle’s departure time, it is transferred to the next BS. Since we do not model the service capacity of a BS, all calls are accepted by the BS and correspondingly, counted.
II-B4 Call duration
Call duration has been traditionally modeled by the exponential random variable [9]. However, more recent field studies show that a better fit for the call duration distribution might be the log-normal distribution [10] or a mixture of two log-normal distributions [11]. The latter can model diurnal (longer) and nocturnal (shorter) calls [12] or a mixture of two different services, for instance.
II-B5 Flow between adjacent cells
The vehicular flows in two adjacent BSs can differ from each other. The recorded number of vehicles in one segment of the highway can be smaller or higher than that of the subsequent segment of the highway. This can be due to arterial roads that merge into the highway, see a schematic representation of such a case in Fig. 1. When the incoming flow in the target cell (BS2 in Fig. 1) is smaller than the outgoing flow from the preceding cell (BS1), we randomly remove vehicles together with their corresponding handover calls from the generated data set. The goal is to reflect the route deviation of the vehicles, namely those that leave the highway, into the network statistics too. When the flow in the target cell is larger than the adjacent, preceding cell, no adjustments are made in the handover statistics generated in the preceding cell. The generation of new calls per each vehicle arriving at the cell is done as explained earlier.
III A highway scenario
Figure 2: A map of the section of US50-E El Dorado County highway used in the study (in gray). The red dots indicate the approximate location of the PeMS detectors on that segment, from Pollock Pines to Lake Tahoe Airport. The detectors are listed in Table I together with the distance between consecutive ones.
We consider a sector of the US50-E highway in El Dorado County, California; specifically, the one shown in Figure 2. The Caltrans detectors, whose measurements are used in the study, are listed in Table I in sequential order from left to right when looking at the map. Their approximate location is marked with red dots in Figure 2. The distance between each two consecutive detectors is indicated in Table I.
The density of the PeMS detectors is high at the fragments of the highway that cross populated areas (such as villages and towns; Pollac Pines and Mevers in this particular scenario) and is much lower on the highway sectors of mountainous regions. Long-range BS typically cover rural areas with sparse population, whereas short-range BSs provide coverage in densely populated (urban) regions. Since the density of the road detectors seems to be guided by similar principles—long-term density of the vehicular traffic—as the distribution of BS— population density and the particularities of the terrain—we choose the range of the BSs to match the distance between sensors. This provides us with a variety of different use cases as the BS range (under otherwise equal conditions) defines the handover rates and cell load. For simplicity, we make the assumption that the vehicular traffic is unidirectional, flowing from Pollac Pines to South Lake Tahoe.
TABLE I: PeMS US50-E Detectors Used in the Study
and the Distance Between Consecutive Ones
BS# PeMS detectors miles
1 Mainline VDS 3086071–50EB JWO Sly Park EB $7.67$
2 Mainline VDS 3086081–50EB at Riverton Barn CCTV $1.62$
3 Mainline VDS 320287–50EB At Ice House $13.27$
4 Mainline VDS 320280–Wrights Lake Rd $13.41$
5 Mainline VDS 317706–Echo Summit $3.89$
6 Mainline VDS 3054051–50EB into Luther 50/89 R. $0.99$
7 Mainline VDS 3410061–50EB JEO Pioneer Trl $3.22$
8 Mainline VDS 317715–F St 1.46
IV Mobile Cellular Traffic Prediction: Definition and Solution
Our goal is to predict the mobile cellular traffic load at each BS based on the historical time-series of their mobile cellular volumes and the road traffic time-series observed in the highway segments covered by the BSs. We focus on a single machine learning model, individually trained (with the BS’s own data) for each BS along the highway since our goal is to validate the usefulness of the concept rather than develop a new method. A basic machine learning model allows to abstract the impact of the model from that of the data set and hence, it is easier to validate.
IV-A Problem Formulation
Mathematically, we denote by $\mathbf{x}^{\tau}\in\mathcal{R}^{p}$ a random variable comprising historical measurements of $p$ metrics relevant to a given cell: mobile traffic volume, and/or road traffic volume and speed measurements. Given a time sequence of $M$ such historical observations $\{\mathbf{x}^{\tau-M+1},\dots,\mathbf{x}^{\tau}\}$, our objective is to learn a prediction model $\mathcal{F}$ that can forecast the future mobile traffic load $\hat{x}$ in the cell during the next time step:
$\hat{x}^{\tau+1}=\mathcal{F}(\mathbf{x}^{\tau-M+1},\dots,\mathbf{x}^{\tau}),$ (1)
so that the prediction error is mininized:
$\text{min}\>L(\hat{x},x),$
where the loss function $L(\mathord{\cdot})$ measures the difference between the estimated $\hat{x}$ and observed $x$ mobile cellular traffic.
IV-B A basic machine learning model
Mobile cellular traffic volumes at BSs experience spatial-temporal dependencies. Since we focus on a short-term forecasting problem at a BS (1), the impact of spatial correlations is confined to the adjacent BSs. We account for this phenomenon by the incoming vehicular flow and speed.
The long short-term memory (LSTM) model [13] has been applied in many time-series prediction tasks including mobile cellular traffic volume forecasting. The model we construct is composed by an LSTM layer, followed by a fully connected feedforward neural network (a multilayer perceptron). This dense layer consists of a single unit and no activation function. It is a linear transformation that maps the final LSTM state into a single target value—the estimated number of calls in a BS.
Mathematically, the model can be represented as two functions $f(g),$ where $g(\cdot)$ is the LSTM learning structure, which transforms the input data into new features. The representation function $f(\cdot)$ maps the learned features into a cellular traffic consumption prediction for the BS under consideration.
V Experiments and results
V-A Setting
V-A1 Input variables values
Real road traffic data from Caltrans detectors of the US50-E highway in El Dorado County, California are used in the experiments. The considered time period is 24-week-long from March 28, 2022 to September 9, 2022 (week 13 to week 24 of 2022). The data comprise the flow and speed from Monday to Friday of each week. There are 288 data points per day due to the 5-minute granularity.
We generate the network statistics following the methodology developed in Section II and input values as follows. New calls are generated with arrival rate $\lambda=1/5,$ one call per 5-minute interval on average. When defining the mean call duration and its variance, we are guided by [12] and [10]. We set the mean call duration to 1 minute for exponentially distributed calls and to 1 and 10 minute for the mixture of two log-normally distributed calls, each with equal weight. The variance of the log-normal calls is 3 times larger than their mean [12]. The range of the BSs is listed in Table I.
V-A2 Data
All BSs experience mostly typical for their highway segment mobile volumes yet these traffic loads are quickly varying in short-term scales and are weekly different. For several BSs there are outlying from the rest weeks.
The road metrics—flow and speed—are inversely associated. In addition, their correlation varies over time.
Flow and the generated number of calls are strongly correlated. This reciprocal association, corresponds to recent [14] and earlier studies [15], which show that there exists a strong correlation between population density and mobile cellular use in urban scenarios. Research on population mobility relies on such a high correlation when using call detail records (CDRs) from mobile operators to understand and model human mobility in urban environments [16]. Furthermore, people’s density is often estimated from mobile phone network data [16], [17], [18]. Under the assumption of connected vehicles and a highway scenario—the setting of our study—the load on the BSs that serve the highway is exclusively generated by the road traffic on that highway (the applications used by the autonomous driving vehicles and their passengers). Naturally, the highest correlation is between the flow of vehicles and the new calls they place on their serving BS and correspondingly, the total number of calls. The linear association between the flow and number of handovers is often of a smaller magnitude because not all vehicles have active calls when crossing the borders between BSs, namely when arriving to a new cell.
The number of calls is inversely related to the speed as whenever the road is congested, the dwell time of the vehicle on the road is increased and consequently the probability of initiating a new call grows because we model the call arrival rate as a constant random variable (a variable with variance zero). This is in line with the common observation that the load on the mobile cellular network increases during traffic jams, although we do not model this phenomenon explicitly.
V-A3 Machine learning model
Our model has a single LSTM layer with 16 cells. In the training phase, we use the root mean square propagation  and mean squared error (MSE) loss for optimizing the models parameters. We set the length of the historical traffic series to six samples (30 min) due to the speed with which changes in the traffic volume occur in 5-minute intervals. The prediction horizon is the next 5-minute interval. The 24 weeks of data are split 12:6:6 into training, validation, and testing. Training data is shuffled.
V-B How much can road traffic data help?
V-B1 Metrics
We measure the forecasting performance with the mean absolute error (MAE), mean absolute percentage error (MAPE), MSE and root mean squared error (RMSE).
TABLE II: Prediction Performance on Calls Datasets and Flow, Speed and Calls Dataset
Calls Flow Speed Calls
BS MAE MSE MAPE RMSE MAE MSE MAPE RMSE
3054051 0.434 0.434 0.439 0.485 0.492 0.503 116.0 118.0 128.3 0.697 0.702 0.709 0.364 0.369 0.377 0.341 0.347 0.354 84.3 90.1 93.1 0.584 0.589 0.595
3086071 0.282 0.283 0.287 0.154 0.155 0.157 820.8 837.2 880.6 0.392 0.393 0.396 0.187 0.187 0.187 0.067 0.067 0.068 506.8 515.6 523.9 0.259 0.259 0.261
3086081 0.394 0.396 0.399 0.317 0.318 0.319 335.2 347.4 365.4 0.563 0.564 0.565 0.339 0.34 0.342 0.237 0.238 0.238 293.2 298.3 305.9 0.487 0.487 0.488
317706 0.155 0.156 0.158 0.058 0.059 0.060 127.5 128.4 137.8 0.241 0.242 0.244 0.129 0.13 0.135 0.038 0.039 0.040 107.4 110.8 112.0 0.196 0.198 0.199
317715 0.321 0.324 0.325 0.242 0.243 0.246 103.5 109.3 117.3 0.492 0.493 0.496 0.295 0.297 0.3 0.208 0.21 0.218 87.4 92.4 92.8 0.456 0.459 0.467
320280 0.080 0.081 0.097 0.013 0.013 0.015 54.0 54.6 57.3 0.112 0.114 0.121 0.065 0.066 0.070 0.008 0.008 0.009 47.3 47.7 49.1 0.090 0.091 0.094
320287 0.066 0.067 0.067 0.009 0.009 0.009 86.7 87.0 88.3 0.094 0.094 0.094 0.051 0.052 0.054 0.005 0.006 0.006 64.6 66.6 67.7 0.073 0.075 0.076
3410061 0.404 0.404 0.407 0.389 0.391 0.397 696.4 725.7 757.9 0.624 0.626 0.63 0.354 0.356 0.358 0.309 0.31 0.316 484.8 494.5 503.2 0.556 0.557 0.562
aThe three values listed per error type are the minimum, median and maximum. The MAPE values are in percentage.
TABLE III: Prediction Performance on New, Handover, and Total Calls Dataset and the Same Dataset but with Flow and Speed Time-Series
New HO Calls Flow Speed New HO Calls
BS MAE MSE MAPE RMSE MAE MSE MAPE RMSE
3054051 0.426 0.428 0.431 0.479 0.498 0.504 101.3 103.1 114.5 0.692 0.706 0.71 0.366 0.369 0.37 0.343 0.35 0.355 83.7 88.2 91.9 0.585 0.591 0.596
3086071 - - - - - - - - - - - - - - - - - - - - - - - -
3086081 0.39 0.39 0.392 0.307 0.307 0.309 326.7 337.1 349.7 0.554 0.554 0.556 0.336 0.339 0.346 0.235 0.235 0.238 294.6 300.3 321.9 0.484 0.485 0.488
317706 0.154 0.155 0.158 0.056 0.057 0.06 128.1 131.7 134.0 0.237 0.24 0.245 0.131 0.131 0.133 0.039 0.04 0.041 106.6 112.0 114.1 0.199 0.2 0.202
317715 0.323 0.326 0.329 0.242 0.245 0.25 106.3 111.3 113.4 0.492 0.495 0.5 0.299 0.304 0.306 0.215 0.218 0.233 82.2 85.5 91.9 0.463 0.467 0.482
320280 0.076 0.081 0.100 0.012 0.012 0.015 54.0 55.5 56.4 0.108 0.111 0.123 0.062 0.063 0.069 0.008 0.008 0.008 44.9 46.8 48.6 0.087 0.089 0.092
320287 0.066 0.066 0.067 0.009 0.009 0.009 87.8 88.7 90.6 0.093 0.094 0.094 0.052 0.053 0.055 0.005 0.006 0.006 65.5 66.0 70.4 0.073 0.075 0.078
3410061 0.405 0.408 0.415 0.394 0.403 0.423 590.1 639.2 655.4 0.628 0.635 0.65 0.355 0.357 0.363 0.312 0.316 0.322 458.7 474.7 483.2 0.558 0.562 0.568
aBS 3086071 is the first one in our scenario. Therefore, there is no handover traffic from a preceding BS.
TABLE IV: Prediction Performance with Gaussian Noise Added to the Flow
Flow Speed Calls Flow Speed New HO Calls
BS MAE MSE MAPE RMSE MAE MSE MAPE RMSE
3054051 0.369 0.371 0.375 0.348 0.354 0.358 84.0 85.9 87.8 0.59 0.595 0.598 0.367 0.368 0.372 0.343 0.348 0.354 86.3 89.4 92.7 0.586 0.589 0.595
3086071 0.205 0.205 0.206 0.077 0.077 0.078 594.5 599.4 602.1 0.278 0.278 0.279 - - - - - - - - - - - -
3086081 0.346 0.349 0.35 0.244 0.245 0.247 293.1 315.5 320.6 0.494 0.495 0.497 0.344 0.347 0.348 0.241 0.242 0.243 311.9 314.9 315.8 0.491 0.492 0.493
317706 0.134 0.136 0.138 0.04 0.042 0.043 110.0 111.3 119.5 0.2 0.205 0.207 0.135 0.135 0.138 0.041 0.042 0.043 107.5 110.8 117.1 0.202 0.204 0.207
317715 0.301 0.303 0.306 0.214 0.214 0.224 88.5 91.0 96.4 0.462 0.463 0.473 0.303 0.304 0.309 0.216 0.22 0.228 89.7 94.0 95.6 0.465 0.469 0.477
320280 0.070 0.072 0.074 0.009 0.009 0.010 50.5 50.9 51.1 0.096 0.097 0.098 0.071 0.076 0.084 0.009 0.010 0.011 49.8 50.0 52.1 0.095 0.099 0.104
320287 0.057 0.057 0.058 0.006 0.006 0.006 71.7 72.9 74.8 0.079 0.080 0.080 0.056 0.057 0.058 0.006 0.006 0.007 72.9 75.0 76.4 0.078 0.079 0.083
3410061 0.36 0.361 0.365 0.313 0.318 0.327 481.2 495.0 514.7 0.56 0.564 0.572 0.361 0.362 0.369 0.32 0.322 0.334 475.7 496.3 510.8 0.566 0.567 0.578
aBS 3086071 is the first one in our scenario. Therefore, there is no handover traffic from a preceding BS.
V-B2 Overall performance
The results from the four different data sets—1)  calls (denoting total number of calls, the sum of new and handover calls) only; 2) flow, speed and calls; 3) new, handover, calls and 4) flow, speed, new, handover, calls—when the call duration is modeled by a mixture of two log-normal distributions are summarized in Table II and Table III. For each error metric we show the minimum, median and maximum of the 5 simulation runs conducted for each data set and BS. We make the observations as follows:
  • Road traffic data consistently improve the mobile cellular traffic prediction performance. All BSs observe improvements in all their forecasting performance measures when learning from both network and road data.
    The prediction error reduction is between $8.4\%$ and $33.9\%$ (MAE), $13.3\%$ and $56.5\%$ (MSE), $12.5\%$ and $38.4\%$ (MAPE), and $6.9\%$ and $34\%$ (RMSE) when in addition to (total number of) calls, also flow and speed are employed, and when comparing the medians of all BSs. Similar percentage decrease is recorded for the minimum and maximum errors. Using more refined network KPIs—those that contain the number of new and handover calls, and their total—and road data, reduces the prediction errors too: between $7.5\%$ and $21.8\%$ (MAE), $11\%$ and $35.4\%$ (MSE), $10.9\%$ and $25.7\%$ (MAPE), and $5.7\%$ and $19.7\%$ (RMSE) for the medians and likewise for the smallest and largest errors.
  • The road traffic metrics can capture the processes underlying the mobile cellular traffic generation. We employ flow and speed as a means to characterize the underlying population dynamics and through them the data generation process. The flow captures the number of potential call generation sources and the speed can serve as an indicator for road traffic congestion and hence, for increased traffic load on the BS. The number of handovers (or equivalently, the handover rate) is an indicator of mobility too yet our results are not conclusive about its impact. Contrasting the performance of calls to that of new, handover, and calls, we observe an error improvement in some of the error measures: it is in the order of no more than $3\%$ for the majority of the BSs.
    Finally, employing flow, speed and calls (3 variables) leads to a large reduction in the prediction errors even when contrasted with employing new, handover and calls (3 variables): $8.6\%$ and $22.8\%$ (MAE), $14.2\%$ and $36.3\%$ (MSE), $11.5\%$ and $25\%$ (MAPE), and $7\%$ and $20.2\%$ (RMSE) for the medians.
  • Choice of data (split) and hyper-parameter tuning.
    We highlight that the hyper-parameter tuning is an indispensable part of every machine learning process. In the tables above we show the results from using 24 weeks of data and 75:25 ratio of train and validation to test data. This amount of data seems to be beneficial mainly to the BSs with long range (320280 and 320287) but less to the reaming BSs, which evidence improved performance when trained over shorter periods. We also observe cases when the performance of some of the BSs is decreased when using road data. This is the case for the MSE and RMSE of BS 320287 when the data set consists of weeks 15 to 21 and the data is chronologically split  into 3:2:2 for training, validation and testing [^2] See https://github.com/nvassileva/DataMatters for detailed results.. We identify week 15 as outlying from the other training weeks and consequently, as the main explanation of the results. For the same period but with 4:1:3 partition of the data, the results are consistent with what we report above—decreased prediction errors across all performance measures and BSs. This 4:1:2 data split also turns into smaller forecasting errors for the purely network data sets. In other words, choice of data that captures the underlying patterns and an appropriate hyper-parameter tuning are always required. In our on-going work we are studying the effect of the data split on the overall forecasting performance.
V-B3 Sensitivity to road metrics accuracy
We envision that road traffic measurements can directly be fed into the learning algorithms as smart cities and roads are equipped with detectors that can provide such critical information. Another alternative is to use road traffic predictions instead of measurements. The machine learning research community has seen a surge of deep learning methods that tackle this prediction task with several solutions that achieve high levels of forecasting accuracy.333 An advantage of such an approach is that the mobile cellular prediction algorithms can receive early information about the vehicular flow and speed that will be evidenced in the BS during the next time slot. This in advance received data can increase further the prediction accuracy of the algorithms. In fact, our preliminary results validate this concept under similar setting as those we reported above and experimented with.
To assess the sensitivity of the mobile cellular predictions to estimation errors in the road predictions, we introduce an estimation error in the flow measurements by the PeMS detectors. In particular, we assume a prediction error of $5\%$ in the number of vehicles per 5-minute intervals. The speed, although time correlated with the flow, especially during rush hours, is not as strongly correlated with the mobile cellular traffic for the majority of the BS. Hence, we do not model speed prediction errors. We add Gaussian noise to the real measurements obtained via PeMS. In particular,
$\hat{v}=v+\epsilon,\qquad\epsilon\sim\mathcal{N}(0,\sigma),$ (2)
where $v$ denotes the real PeMS flow data and $\hat{v}$ is the estimated flow. Whereas our learning model is fed with the estimated value of the flow variable, we generated the mobile cellular traffic with the true PeMS measurement.
The results are reported in Table IV and show that although the noise in the flow variable decreases cell load prediction accuracy, employing road data is largely beneficial yet. Overall, across all error measures and all BSs and the 4 data sets, the prediction error improvement is between $5.2\%$ and $49.9\%$ for the median when using estimated (with errors) road traffic time series data.
Employing flow, speed and calls—when flow is reported with errors—leads to a reduction in the prediction errors even when contrasted with using new, handover and calls data set: the improvement is between $6.4\%$ and $29\%$ for the medians among the 4 error measures we use.
VI Conclusion
In the past, mobile cellular data have been used to model population dynamics and to enrich our understanding of urban mobility. We take a different look and use population dynamics in a highway context to forecast mobile cellular traffic load. By employing vehicular flow and speed time-series observed in the base stations serving a highway, the traffic load prediction error at each base station can be largely and consistently reduced: between about $5\%$ and $56.5\%$ depending on the error measure, base station and specific data sets. This is a powerful result that can be employed in deep learning models such as graph neural networks for mid- to long-term cellular load forecasting or federated learning for short-term predictions.